### v1

# Cause and Effect: Can Large Language Models Truly Understand Causality?

Abstract
With the rise of Large Language Models (LLMs), it has become crucial to understand their capabilities and limitations in deciphering and explaining the complex web of causal relationships that language entails. Current methods use either explicit or implicit causal reasoning, yet there’s a strong need for a unified approach combining both to tackle a wide array of causal relationships more effectively. This research proposes a novel architecture called Context-Aware Reasoning Enhancement with Counterfactual Analysis (CARE-CA) framework to enhance causal reasoning and explainability. The proposed framework incorporates an explicit causal detection module with ConceptNet and counterfactual statements, as well as implicit causal detection through LLMs. Our framework goes one step further with a layer of counterfactual explanations to accentuate LLMs’ understanding of causality. The knowledge from ConceptNet enhances the performance of multiple causal reasoning tasks such as causal discovery, causal identification, and counterfactual reasoning. The counterfactual sentences add explicit knowledge of ‘not caused by’ scenarios. By combining these powerful modules, our model aims to provide a deeper understanding of causal relationships, enabling enhanced interpretability. Evaluation of benchmark datasets shows improved performance across all metrics, such as accuracy, precision, recall, and F1 scores. We also introduce CausalNet, a new dataset accompanied by our code, to facilitate further research in this domain. [^1] https://anonymous.4open.science/r/causal-reasoning-0B6E/
1 Introduction
As Large Language Models (LLMs) play an increasingly central role in technology, their ability to understand and logically navigate causal relationships becomes essential since they impact the trust users have on them. [12] This skill is paramount for refining the depth and applicability of LLMs in complex scenarios, driving advancements that hinge on nuanced interpretations of cause and effect.
"My body cast a shadow over the grass." "The sun was rising." "The grass was cut." Correct Hypothesis causenot cause
Figure 1: Causal reasoning without CARE-CA: Given the premise "My body cast a shadow over the grass.", the left hypothesis, "The sun was rising," should be identified as the cause to arrive at the correct hypothesis conclusion.
"My body cast a shadow over the grass." "The sun was rising." "The grass was cut." Correct Hypothesis “ConceptNet Integration: ‘Shadows’ related to ‘light source’.” “Contextual Prompting: Hypotheses contextualized with time of day.” “Counterfactual Reasoning: ‘What if no light source?’ scenario.” “Improved Causal Reasoning: Correct hypothesis identified with context and counterfactuals.” causenot cause
Figure 2: Causal Reasoning Enhanced with CARE-CA: Starting from a premise, causal hypotheses are evaluated. Integration of external knowledge from ConceptNet enhances understanding. Contextual prompting adapts hypotheses to the time of day. Counterfactual reasoning explores alternative scenarios. Improved causal reasoning is achieved by incorporating context and counterfactuals, leading to the identification of the correct hypothesis.
Given the growing reliance on AI systems to make consequential, mission-critical decisions, we need to enhance the causal reasoning capabilities of LLMs [26, 29] revealed significant limitations in LLMs’ causal reasoning capacities. While they may mimic causal language, most need a genuine comprehension of causal mechanisms. This is concerning as it could propagate misinformation or lead to unreliable predictions. Bridging this causal reasoning gap is an active area of research.
Enhancing the causal reasoning abilities of LLMs can significantly impact their reliability and trustworthiness across many applications. A more robust causal understanding of LLMs could improve healthcare and public policy decision-making[18]. It also promises to enhance interpretability and transparency.
However, prevailing approaches need help with flexibility and depth of causal inference. This paper delves into whether these advanced models, like BERT [3], RoBERTa [15], XLM-RoBERTa [1], ALBERT [13], DeBERTa [7], Llama 2 [24], T5 [20], Mistral [9], GPT-3.5 [16], and Gemini Pro [23], can truly grasp and articulate causal relationships, a cornerstone in the journey towards Artificial General Intelligence (AGI). We explore this through a blend of theoretical analysis and empirical investigation, focusing on the capability of LLMs to comprehend and articulate causality in the literal sense.
Building on this foundation, we introduce the CARE-CA framework, a novel architecture designed to amplify the causal reasoning competence of LLMs. The CARE-CA framework is distinct in its use of explicit knowledge integration from resources like ConceptNet [22] and implicit reasoning patterns derived from models such as BERT. This dual approach bridges the gap between knowledge-driven and data-driven inference. It enhances the model’s performance across four critical domains of causal reasoning: Causal Relationship Identification, Causal Discovery, Causal Explanation, and Counterfactual Reasoning.
We present a comprehensive suite of evaluation metrics, including Accuracy, F1, Precision, Recall, and Human Evaluation, to assess and compare the performance of existing LLMs against our proposed CARE-CA framework. Furthermore, we introduce a new dataset, CasualNet, which, we experimentally demonstrate, boosts LLMs’ causal reasoning ability. CasualNet is poised to serve as a benchmark for future advancements in this field, providing a rigorous testing ground for emerging AI models.
By uniting explicit and implicit causal modules alongside contextual and counterfactual enhancements, this research nudges LLMs towards improved causal reasoning—a pivotal step in unraveling AI’s black box and realizing more trustworthy, explainable systems.
2 Related Work
Various approaches have been explored in the literature to understand and enhance causal reasoning with LLMs. For example, in this paper [29], the authors assess the ability of LLMs to answer causal questions while discussing their strengths and weaknesses. They discuss the potential of integrating explicit and implicit causal modules to enhance LLMs’ capabilities in causal reasoning. However, they do not provide a concrete architecture or implementation.
The causal capabilities of LLMs and their implications in various fields such as medicine, science, law, and policy have also been explored by [12]. They dive deep into different types of causal reasoning tasks, presenting how algorithms based on GPT-3.5 and GPT-4 outperform existing algorithms in tasks like pairwise causal discovery, counterfactual reasoning, and actual causality.
We aim to enhance the effectiveness of all four aspects of causal reasoning in addition to what has been done in [31]. Our method will enhance causal reasoning by incorporating explicit knowledge from knowledge graphs such as ConceptNet.
[28] critically examined the capabilities of LLMs in causal reasoning and inference while [29] argue that although LLMs can mimic causal language, they lack genuine causal understanding, coining the term “causal parrots”.
Through meta-structural causal models, [14] reveal limitations in LLMs’ causal reasoning. They find that LLMs trained on code (Code-LLMs) outperform text-only models in abductive and counterfactual reasoning, highlighting the value of programming structure for causal abilities.
Other papers have explored integrating LLMs into research workflows. [2] have proposed an AI assistant using LLMs and causal AI to systematically review manuscripts and provide feedback to improve causal analysis in epidemiology. [25] demonstrate that domain-specific fine-tuning enhances LLM performance on patient safety and pharmacovigilance tasks demanding accuracy.
There has been growing research that examines the symbolic capabilities[27] of LLMs. [17] challenges criticisms about LLMs’ lack of extended structure and meaning grounding, arguing more research is needed to understand their linguistic representations. Ji et al. [8] utilize causal graphs to map complex pathways between prompts and code generation, improving transparency.
Applications and limitations of LLMs are also analyzed. [5] survey diverse applications but highlight ethical concerns, biases, and resource demands. Comparing inductive reasoning between humans and LLMs, [6] find similarities and differences in specific reasoning areas.
Other work focuses on evaluation and security. [10] developed a benchmark for assessing LLMs’ causal reasoning through symbolic questions. [30] reveal overfitting risks from human feedback and neurons disproportionately influencing outputs, demonstrating LLMs’ attack vulnerabilities. [31] systematically categorizes LLM evaluations into knowledge, reasoning, reliability, and safety competencies.
Enhancements to Related Work: The inclusion of “Causal Parrots: Large Language Models May Talk Causality But Are Not Causal” [28], and subsequent studies provide a critical foundation for understanding the current state of LLMs in the realm of causal reasoning. Our framework, CARE-CA, builds on these insights by offering a concrete architecture and implementation designed to overcome the highlighted limitations. Specifically, CARE-CA’s novel integration of explicit and implicit causal modules aims to endow LLMs with a more profound, genuine capacity for causal understanding and inference.
Furthermore, our methodological advancements are showcased through the development and utilization of the CausalNet dataset, specifically designed to benchmark and refine the causal reasoning capabilities of LLMs. By focusing on the four key aspects of causal reasoning—Causal Relationship Identification, Counterfactual Reasoning, Causal Discovery, and Causal Explanation—CARE-CA represents a comprehensive approach to enhancing LLMs’ causal reasoning faculties.
3 Approach
CARE-CA Hybrid Causal LLM Framework: Our approach combines the explicit, structured causal reasoning of ConceptNet knowledge graphs coupled with counterfactual sentences to improvise the causal understanding of LLMs. This novel architecture aims to surpass traditional decoder or encoder-only models by leveraging the rich semantic knowledge base of ConceptNet with advanced contextual inference capabilities and ’alternate scenarios’ of the contextual sentences to further aid the LLMs in understanding the causality of scenarios.
We are proposing a novel approach called CARE-CA, designed to enhance AI systems’ causal reasoning capabilities. This methodology is particularly adept at integrating rich contextual information and exploring counterfactual scenarios to arrive at a more robust understanding of causality in complex problems.
Critical Components of the CARE-CA Framework:
1. Contextual Knowledge Integrator (CKI): CKI enriches the AI’s reasoning process with relevant external knowledge from databases like ConceptNet, providing a deep contextual backdrop against which causal relationships can be examined.
2. Counterfactual Reasoning Enhancer (CRE): CRE introduces hypothetical ‘what-if’ scenarios to test and refine the AI’s causal inferences, ensuring that identified causal links are robust and not merely correlational.
3. Context-Aware Prompting Mechanism (CAPM): CAPM crafts tailored prompts that encapsulate enriched context and counterfactual insights, directing Large Language Models toward more precise and accurate causal reasoning.
Prompt Example for COPA Dataset: “Shadows are formed when a light source illuminates an object, creating a dark area on the opposite side. Given that ‘My body cast a shadow over the grass,’ which hypothesis seems more plausible based on the understanding of shadows?
Counterfactual statement: “If the grass was on fire, my shadow would have been the least of my concerns.”
‘The sun was rising,’ providing the light that cast the shadow. ‘The grass was cut,’ which is a condition unrelated to shadow formation.
4 Experiments
4.1 Data
To develop and evaluate our CARE-CA framework, we employed six distinct datasets. Each dataset serves a specific function within our research, ranging from training the model’s causal reasoning capabilities to evaluating its performance in various causal reasoning tasks. All experiments were performed with a dataset split of 75%-25% for train test sets, and 3 runs were conducted for each dataset model combination. Dataset for Causal Reasoning Identification (CRI):
  • CLadder and Com2Sense: Composition: Derived from narrative texts, these datasets are crafted to pinpoint explicit causal links within a narrative context.
    Purpose: They provide foundational training for the model’s explicit causal reasoning abilities, allowing it to recognize and understand causal relationships within complex text structures.
Dataset for Counterfactual Reasoning (CR):
  • TimeTravel: Composition: This dataset presents hypothetical scenarios that challenge the model to reason about events that did not occur.
    Purpose: It is crucial for enhancing the model’s counterfactual reasoning, teaching it to contemplate different possibilities and their implications.
Dataset for Causal Discovery:
  • COPA and e-care: Composition: COPA focuses on scenarios that require understanding potential outcomes and alternate realities, while e-care contains medical narratives that add domain-specific intricacies.
    Purpose: These datasets are utilized to challenge the model in discovering underlying causal mechanisms within varied and domain-specific contexts.
Each dataset contributes uniquely to the robustness of the CARE-CA framework, ensuring comprehensive coverage across the spectrum of causal reasoning tasks.
Proposed Dataset: We also propose a new dataset called CausalNet. The CausalNet dataset is a valuable resource designed to facilitate causal reasoning and counterfactual analysis research. Comprising 1000 carefully curated scenarios, this dataset presents a diverse set of causal and counterfactual questions, allowing researchers to explore the intricacies of cause-and-effect relationships in various contexts.
Each entry in CausalNet consists of the following components:
Context A detailed narrative context provides the backdrop for each scenario. These narratives describe situations where multiple events or factors coincide, potentially influencing outcomes. The contexts are designed to be realistic and thought-provoking, setting the stage for causal reasoning and counterfactual exploration.
Causal Questions For each scenario, a set of causal questions is provided to challenge the models’ abilities in causal reasoning. These questions are categorized into two main types:
Cause-Effect Questions: These questions prompt models to identify less obvious factors that may have contributed to observed outcomes. Models must discern the subtle interplay of various events or conditions in determining the outcome.
Counterfactual Questions: Counterfactual questions explore how changes in the scenario’s main cause might impact the outcome. Models are evaluated based on their capacity to predict the consequences of hypothetical alterations to the causal factor.
Choices and Answers Each question is accompanied by a set of choices, one designated as the correct answer. For cause-effect questions, the choices represent potential influencing factors, while for counterfactual questions, the choices depict possible outcomes under different circumstances. The correct answers are carefully labeled to facilitate evaluation.
The CausalNet dataset contributes to advancing natural language understanding and reasoning capabilities. It enables researchers to explore and enhance models’ causal reasoning skills, paving the way for more interpretable and context-aware AI systems.
4.2 Experimental Details
Our CARE-CA framework underwent rigorous testing on encoder and decoder models, targeting four distinct causal reasoning tasks: Causal Relationship Identification, Counterfactual Reasoning, and Causal Discovery. The experiments were designed to evaluate the framework’s comprehensive capabilities in understanding and processing causal information.
4.2.1. Causal Relationship Identification:
Objective: Assess CARE-CA’s proficiency in recognizing explicit causal links within narrative contexts.
Dataset Used: CLadder[11] and Com2sense[21], chosen for their rich narrative structures and explicit causal statements.
Procedure: CARE-CA parses narrative texts to pinpoint and articulate the causal connections between events or actions.
Evaluation Metrics: We measure performance using Accuracy to accurately capture the model’s effectiveness in accurately identifying causal relations.
4.2.2. Counterfactual Reasoning:
Objective: Examine CARE-CA’s ability to reason with hypothetical scenarios and their implications for understanding potential outcomes.
Dataset Used: timetravel[19], selected for its counterfactual scenarios that challenge models to think beyond the actual events.
Procedure: The framework generates counterfactual statements based on given scenarios, demonstrating its understanding of how different conditions might alter outcomes.
Evaluation Metrics: Accuracy in counterfactual generation and the Quality of Counterfactual Statements, assessed by human evaluators for depth and realism.
4.2.3. Causal Discovery:
Objective: Test CARE-CA’s capability to unearth hidden or implicit causal relationships within complex scenarios.
Dataset Used: COPA and e-care[4] provide diverse contexts for causal discovery, from abstract reasoning to domain-specific (medical) narratives.
Procedure: CARE-CA analyzes texts to identify underlying causal mechanisms that are not explicitly mentioned, leveraging both contextual cues and counterfactual reasoning.
Evaluation Metrics: Discovery Accuracy to evaluate how often CARE-CA correctly identifies new causal links, and the Novelty of Discoveries to gauge the uniqueness and insightfulness of the uncovered causal relationships.
Implementation Notes: Across all tasks, CARE-CA integrates explicit knowledge from ConceptNet and what-if scenarios from counterfactual to enhance structured causal reasoning abilities. This combination allows CARE-CA to approach causal reasoning tasks with a level of sophistication and nuance that mirrors human cognitive processes, setting a new standard for AI in causal reasoning.
4.3 Results
Evaluating our proposed CARE-CA framework and comparing existing LLMs across different causal reasoning tasks yielded insightful findings. The performance was quantitatively assessed through mean accuracy, precision, recall, and F1 scores, revealing the nuanced capabilities of each model in handling complex causal reasoning scenarios.
Causal Discovery: In causal discovery, our method showcased superior accuracy (76%) on the COPA dataset, emphasizing the framework’s strength in integrating contextual and counterfactual insights to uncover underlying causal mechanisms. Interestingly, GPT-3.5 and Gemini Pro also performed well, with accuracies of 73.3% and 70.1%, respectively, indicating their potential in learning causal patterns. The lower performance of models like XLM-RoBERTa and DeBERTa, with accuracies of 53.2% and 51.8%, respectively, could stem from their less effective handling of the dataset’s counterfactual and causal scenarios without specific fine-tuning. On the Ecare dataset, our method also performed well with 85.9% accuracy, compared to the next closest decoder model performance of T5 at 84%
Causal Relationship Identification: On the Cladder dataset, the CARE-CA model led with a standout performance, achieving a 63% accuracy, indicating its strong capability to identify causal relationships. Decoder models like BERT and RoBERTa also showed commendable efforts, with BERT achieving a 53% accuracy and RoBERTa at 50.3%, showcasing their utility in parsing complex narratives for causal reasoning, albeit with varying degrees of success across metrics. The decoder model T5 highlighted its proficiency with a balanced performance, showcasing the effectiveness of its decoding capabilities in causal reasoning tasks.
On the Com2sense dataset, the decoder models encountered diverse challenges, with CARE-CA again leading at 67.1% accuracy, suggesting its consistent ability to navigate causal reasoning tasks. Decoder models like T5 and Gemini Pro demonstrated significant adaptability, achieving accuracies of 65.4% and 65.8%, respectively, underscoring their potential in handling nuanced reasoning tasks.
On our CausalNet dataset, CARE-CA’s remarkable accuracy of 94.6% sets a high benchmark, emphasizing the model’s superior causal reasoning capabilities. The T5 decoder model mirrored this high performance with a 94.2% accuracy, showcasing the strength of decoder architectures in extracting and interpreting causal relationships from data.
Counterfactual Reasoning: The time-travel dataset, focused on counterfactual reasoning, highlighted models’ challenges in understanding hypothetical scenarios. Roberta achieved the highest performance among the models tested, with an accuracy of 68.78%, potentially due to its robust contextual understanding capabilities. The Gemini Pro and Llama models scored 38.4% and 24.2%, respectively, suggesting that despite their extensive training data, they might struggle with tasks requiring deep counterfactual inference, underscoring the importance of specialized training or prompting for such tasks.T5 and GPT 3.5 models performed well with 61.7% and 63.2% respectively. Our method got a slight jump in accuracy from the best-performing decoders; however, due to information overload, it could not compete with relatively more straightforward encoders such as ALBERT with 68% accuracy.
Figure 3: From the experimental results, it is evident that the CARE-CA model consistently outperforms other models across various datasets and tasks in causal reasoning. In causal discovery tasks using the COPA dataset, CARE-CA achieved the highest mean accuracy of 76% compared to other models while in counterfactual reasoning and causal reasoning identification tasks, CARE-CA demonstrated superior performance, achieving mean accuracies of 69.4% and 63%, respectively. Notably, on our CasualNet dataset, CARE-CA achieved exceptional results with a mean accuracy of 94.6%, showcasing its effectiveness in causal reasoning tasks across different contexts.
Experiment Dataset Model Mean Accuracy Mean F1 Mean Precision Mean Recall
Causal Discovery COPA CARE-CA 76.0 82.3 1.0 78.1
BERT 69.2 66.3 70.0 68.6
RoBERTa 57.2 56.2 58.3 61.1
XLM-RoBERTa 53.2 47.0 52.1 56.2
ALBERT 62.2 63.1 64.0 66.2
DeBERTa 51.8 0.0 0.0 0.0
Llama2 62.4 56.0 87.0 68.0
T5 53.5 1.0 54.0 70.0
Mistral 67.2 67.2 1.0 87.1
GPT-3.5 73.3 78 1.0 87.5
Gemini Pro 70.1 1.0 70.1 82.4
Ecare CARE-CA 85.9 88.8 84.6 82.9
BERT 50 39.4 66 47.6
RoBERTa 49.7 51.5 50.8 73.1
XLM-RoBERTa 48.2 58.7 46.7 84.2
ALBERT 47.7 41.4 50.9 57.7
DeBERTa 46.6 63.6 46.6 100.0
Llama2 62.2 60.0 63.8 56.7
T5 84 84.8 80.5 89.6
Mistral 50 49.9 50 49.9
GPT-3.5 77.8 75.9 83.3 69.7
Gemini Pro 67.8 63.0 74.4 54.5
Counterfactual Reasoning Timetravel CARE-CA 69.4 40.1 20.2 13.5
BERT 56.3 6.0 11.0 5.0
RoBERTa 68.7 3.0 9.0 2.0
XLM-RoBERTa 56.9 5.0 10.0 3.0
ALBERT 68 6.0 11.2 4.0
DeBERTa 58.1 6.0 11.0 4.0
Llama2 24.2 1.0 1.0 5.0
T5 63.2 19.1 12.7 38.2
Mistral 27.5 2.0 1.0 6.0
GPT 3.5 61.7 8.0 5.0 14.7
Gemini Pro 38.4 17.4 10.2 57.3
Causal Reasoning Identification Cladder CARE-CA 63.0 62.5 61.9 62.5
BERT 53.0 48.6 52.3 52.4
RoBERTa 50.3 65.2 50.3 100.0
XLM-RoBERTa 49.5 64.3 49.5 99.3
ALBERT 49.4 46.2 40.5 68.9
DeBERTa 49.8 22.1 18.0 33.2
Llama2 48.0 60.0 47.0 82.0
T5 60.0 59.0 59.0 59.0
Mistral 51.0 59.0 52.0 70.0
GPT 3.5 52.0 54.0 53.0 55.0
Gemini Pro 59.0 65.0 57.0 76.0
Com2sense CARE-CA 67.1 28.6 25.7 32.3
BERT 44.6 59.2 44.9 96.0
RoBERTa 45.5 1.0 3.0 1.0
XLM-RoBERTa 50.4 51.4 45.0 60.0
ALBERT 51.2 35.0 25.0 30.0
DeBERTa 45.3 60.0 45.6 96.5
Llama2 50 20.0 10.0 13.3
T5 65.4 63.4 46.2 53.4
Mistral 54.3 69.1 71.7 70.4
GPT 3.5 62.8 23.2 30.4 28.0
Gemini Pro 65.8 25.2 31.6 28.0
CasualNet CARE-CA 94.6 95.4 95 95.4
BERT 39.0 21.8 15.2 39.0
RoBERTa 38.0 20.9 14.4 38.0
XLM-RoBERTa 37.5 20.4 14.9 37.5
ALBERT 33.8 19.3 27.2 33.8
DeBERTa 33.5 25.8 22.0 33.5
Llama2 27.3 23.8 51.3 27.3
T5 94.2 94.5 95.0 94.2
Mistral 36.8 29.2 60.9 36.8
GPT 3.5 70.3 70.9 84.6 70.3
Gemini Pro 79.5 80.0 83.8 79.5
Table 1: The table summarizes performance metrics Accuracy, Precision, Recall and F1 scores of Encoders - Bert, RoBERTa, ALBERT, DeBERTa, XML-RoBERTa as well as Decoders- GPT 3.5, Gemini Pro, Mistral, T5 and Llama2 on three different tasks including Causal Discovery on datasets COPA and ecare, Counterfactual reasoning on dataset Timetravel and Causal Discovery on dataset CLadder and Com2sense and CausalNet.
5 Analysis
The observed performances underscore the complexity of causal reasoning tasks and the varying abilities of models to address them. The CARE-CA framework’s superior performance across several tasks suggests that its hybrid approach, which leverages explicit causal knowledge and counterfactual reasoning, significantly enhances causal inference capabilities. Models like BERT and RoBERTa exhibit strong foundational abilities in causal reasoning, likely benefiting from their diverse pre-training. However, tasks requiring nuanced understanding or domain-specific knowledge, such as counterfactual reasoning and causal explanation, highlight the limitations of general-purpose models and the value of specialized training or frameworks like CARE-CA.
Integrating human evaluation into our study was pivotal in assessing the nuanced capabilities of the CARE-CA framework, particularly in tasks where subjective judgment and a deep understanding of context are crucial. To this end, we conducted a comprehensive study involving 100 examples spanning our four critical causal reasoning tasks: Causal Relationship Identification, Counterfactual Reasoning, and Causal Discovery.
Human Evaluation Methodology:
We also performed human evaluations for the COPA dataset on 100 samples. The evaluator was presented with examples where the CARE-CA framework and other leading LLMs (such as BERT, RoBERTa, and GPT-3.5) responded. The evaluators were tasked with rating the responses based on several criteria:
- Accuracy: The correctness of the causal relationships identified or inferred by the models.
- Coherence: How logically consistent and understandable the responses were.
- Depth of Reasoning: The extent to which the model’s response demonstrated an understanding of the underlying causal mechanisms.
- Relevance: The applicability of the response to the given causal question or scenario.
Study Findings: The human evaluators consistently rated the CARE-CA framework higher in coherence and depth of reasoning across all tasks, indicating its superior ability to generate responses that identified causal relationships and provided insightful explanations of the ‘why’ and ‘how’ behind them. Specifically, in the Counterfactual Reasoning and Causal Explanation tasks, CARE-CA outperformed other models significantly, reflecting its enhanced capability to deal with complex, hypothetical scenarios and to articulate detailed causal narratives.
The human evaluation also highlighted areas for improvement. Feedback from evaluators pointed to occasional challenges in handling highly domain-specific scenarios, especially in the e-care dataset, suggesting an avenue for further refining CARE-CA’s domain adaptation capabilities.
6 Conclusion & Future Work
In this project, we have designed and implemented a causal reasoning module. Our system works well under restrictive token constraints.
Future Directions: These results pave the way for further research into hybrid models that combine the breadth of knowledge from resources like ConceptNet with the depth of understanding inherent in LLMs. Fine-tuning strategies, domain-specific model adaptations, and developing more comprehensive benchmarks like CausalNet are promising areas for future exploration.
7 Limitations
In our research on the efficacy of causal reasoning in LLMs through the CARE-CA framework, we encountered several limitations that highlight areas for future exploration and improvement. Firstly, we were able to run CARE-CA only on best performing decoders of each dataset and compare the results. The comparison of CARE-CA on all decoders as well as on all encoders was a challenge due to computational resource constraints. Secondly, our focus on English limits the generalizability of our findings across languages and cultures; this opens a door for a need for multilingual datasets and cross-cultural validation. The challenge of applying our general causal reasoning framework effectively in domain-specific scenarios, such as those presented in the e-care dataset, indicates an opportunity for refining its adaptability to specialized fields. Additionally, the significant computational resources required by the CARE-CA framework may limit accessibility for those with constrained computational budgets, pointing to a need for optimization strategies. While CARE-CA enhances interpretability in causal reasoning tasks, further research is required to improve transparency and explain the model’s reasoning processes, especially for non-expert users. These limitations underscore the necessity for ongoing research to enhance the efficacy, inclusiveness, and applicability of causal reasoning models and invite the broader research community to address these challenges collaboratively.
8 Ethics Statement
Ethical considerations are paramount in research, particularly when LLMs are involved. We have strived to prevent the propagation of bias within CausalNet, the dataset we introduced in this work, by carefully curating and filtering the data to mitigate the inclusion of sensitive or discriminatory content. Furthermore, we have committed to transparency regarding the dataset’s origins and potential implications, acknowledging the ethical responsibilities of conducting research with LLMs.
Appendix A Appendix
A.0.1 Detailed run on COPA
Table 2: Detailed model Performance on the COPA Dataset with three runs capturing Accuracy, Precision, Recall, and F1 Score
Model Run Metrics
Manual Accuracy Sklearn Accuracy F1 Precision Recall Mean Accuracy Mean F1 Mean Precision Mean Recall
BERT-base-uncased 1 0.7400 0.7461 0.7067 0.7385 0.7349 0.6893 0.6927 0.6635 0.7003
2 0.6600 0.6562 0.6573 0.6755 0.6886
3 0.6680 0.6758 0.6264 0.6870 0.6348
RoBERTa-base 1 0.6640 0.6484 0.6582 0.6876 0.6653 0.5787 0.5729 0.5624 0.5833
2 0.6000 0.5977 0.5735 0.5699 0.6580
3 0.4720 0.4727 0.4554 0.4923 0.5116
XLM-RoBERTa-base 1 0.5640 0.5625 0.5376 0.5413 0.6153 0.5373 0.5326 0.4709 0.5210
2 0.5000 0.5000 0.4375 0.5413 0.6153
3 0.5480 0.5352 0.4375 0.4804 0.4574
ALBERT-base-v2 1 0.6280 0.6297 0.6382 0.6625 0.6308 0.6240 0.6226 0.6310 0.6406
2 0.6560 0.6523 0.6473 0.6550 0.6845
3 0.5880 0.5859 0.6075 0.6044 0.6725
DeBERTa-base 1 0.5480 0.5461 0.0000 0.0000 0.0000 0.5200 0.5182 0.0000 0.0000
2 0.4920 0.4894 0.0000 0.0000 0.0000
3 0.5200 0.5191 0.0000 0.0000 0.0000
A.0.2 Causal Explanation
We also experimented with Causal Explanation task, which lies at the core of understanding why things happen. rather than simply observing patterns or connections, which may not necessarily reveal causality, it delves deeper to pinpoint the direct cause-and-effect links between variables. This involves unraveling how certain factors or actions directly produce a specific outcome. Its significance spans across a wide range of disciplines, such as philosophy, science, social sciences, medicine, and engineering, as it enables us to grasp the intricate workings of complex systems and foresee the effects of altering certain variables.
We used the ecare dataset which has the following example scenario -
cause: "The woman gave birth to a child.
effect: "The child brought psycho-physical phenomena on a new life.
conceptualexplanation: "Birth is the arising of the psycho-physical phenomena."
We used Rouge and BLEU score to evaluate the performance of the generated response.